### v2

# LLM-Based Agentic Systems for Software Engineering: Challenges and Opportunities

yongjian.tang@siemens.com0009-0008-8670-6843 thomas.runkler@siemens.com0000-0002-5465-198X
LLM-Based Agentic Systems for Software Engineering: Challenges and Opportunities
Yongjian Tang Siemens AG
Garching near Munich
Germany
Technical University of Munich
Garching near Munich
Germany
Thomas Runkler Siemens AG
Garching near Munich
Germany
Technical University of Munich
Garching near Munich
Germany
Zusammenfassung
Despite recent advancements in Large Language Models (LLMs), complex Software Engineering (SE) tasks require more collaborative and specialized approaches. This concept paper systematically reviews the emerging paradigm of LLM-based multi-agent systems, examining their applications across the Software Development Life Cycle (SDLC), from requirements engineering and code generation to static code checking, testing, and debugging. We delve into a wide range of topics such as language model selection, SE evaluation benchmarks, state-of-the-art agentic frameworks and communication protocols. Furthermore, we identify key challenges and outline future research opportunities, with a focus on multi-agent orchestration, human-agent coordination, computational cost optimization, and effective data collection. This work aims to provide researchers and practitioners with valuable insights into the current forefront landscape of agentic systems within the software engineering domain.
keywords:
LLMs Agents Software Engineering Future Challenges
1 Introduction
Large Language Models (LLMs) [64] have shown promising potential to transform the Software Development Lifecycle (SDLC) . Their exceptional capabilities in text understanding, knowledge inference, and reasoning have opened new avenues for addressing complex software engineering tasks. Both academia and industry have explored a wide range of techniques to address various Software Engineering (SE) use cases, such as prompt engineering [39, 45], in-context learning [9, 46, 47], parameter-efficient fine-tuning [16], and retrieval augmented generation [14]. Each approach offers distinct advantages for tailoring LLM capabilities to specific SE tasks.
Building on these foundational techniques, LLM-based Multi-agent systems have emerged as a promising paradigm to address the growing complexity of modern software development. These systems can orchestrate multiple specialized agents, integrate diverse development tools, access external knowledge bases, and leverage domain-specific resources to provide comprehensive, end-to-end solutions for real-world SE challenges.
This work is structured as follows: Section 2 introduces the backgrounds and foundations of language models and agentic systems. Section 3 presents a systematic literature review and investigates existing multi-agent approaches across requirements engineering, code generation, static code checking, testing, and debugging. Section 4 outlines our methodological approaches and research agenda, covering cutting-edge language models, SE benchmarks, and agentic frameworks. Section 5 proposes future challenges and opportunities, with a focus on agent capability enhancement, human-agent coordination, data collection, and cost optimization.
2 Backgrounds and Foundations
2.1 Transformers and Language Models
With the advent of transformer architecture [48], language modeling has received significant attention. The highly parallelizable architecture not only scales well with massive data and computing power but also effectively captures long-range dependencies and contextual relationships in text. BERT [8] initially stands out among all models. The semantic representations obtained in pre-training make it approachable for a broad spectrum of downstream tasks in specific domains. This ”pre-training and fine-tuning” mode has inspired numerous follow-up works. Meanwhile, the research community continues to improve the performance of language models by scaling up their sizes. Compared with smaller counterparts, large-scale models demonstrate unseen emergent abilities [52] and have provided us with more problem-solving possibilities, such as prompt engineering, in-context learning, and step-by-step reasoning.
2.2 LLM-based Agentic Systems
Facilitated by these LLMs, traditional agent-based systems evolved significantly. Such LLM-based agentic systems are now capable of handling fuzzy textual input, learning from demonstration examples, generating more contextually relevant responses, thus providing more coherent and meaningful interaction with users.
The main advantages of such collaborative agentic systems can be summarized as follows:
  • Each agent can be specialized for its specific task. For instance, one agent can focus on code generation, another on testing, another on documentation, etc. Specialization allows each agent to be optimized for its specific task, leading to superior performance and efficiency.
  • Multi-agent systems are inherently modular. If a particular agent needs a new capability, it can be upgraded independently through fine-tuning or prompting without affecting the entire system. This modularity facilitates better maintenance and easier integration of new technologies.
  • Multi-agent systems benefit from collaborative problem-solving capabilities, with each agent contributing a unique perspective, resulting in more versatile solutions.
  • Multi-agent systems excel in using tools and other resources, leading to more efficient and effective outcomes than those achieved by a single large language model.
  • Multi-agent systems can perform tasks in parallel, speeding up the software development.
3 Literature Review – Agents for SE
A wide range of studies have explored the integration of LLM-based multi-agent systems across different stages of SDLC. We present a concise review of selected works, categorizing them into five pivotal fields of software engineering – requirements engineering, code generation, static code checking, testing, and debugging.
3.1 Requirements Engineering
Requirements engineering focuses on defining and managing system requirements. This stage consists of several interconnected processes: elicitation, modeling, specification, analysis, and validation. These processes work collectively to ensure that requirements align with stakeholders’ needs while adhering to pre-defined quality standards.
Many LLM-based agents have been proposed to address the challenges in this stage. For instance, a multi-agent system [40] is built to generate and prioritize use stories from the initial project description, effectively identifying the core features of the application and thus enhancing the requirement elicitation and analysis. Furthermore, a framework called SpecGen [32] divides specification generation into two phases. The first phase employs a conversational approach, guiding the LLM to generate relevant specifications for a given program. The second phase applies four mutation operators to the model-generated specifications and selects verifiable specifications from the mutated ones through a heuristic selection strategy. MARE [22], another multi-agent system covering several phases of requirements engineering, performs nine actions to generate requirement models and specifications, surpassing SOTA baselines using text parser and logical reasoning by 15.4%.
3.2 Code Generation
Code generation has been a long-standing focus of SE research. A prominent multi-agent setup for coding typically relies on role specialization and iterative feedback loops to optimize collaboration among agents. Common roles include orchestrator, programmer, reviewer, tester, and information retriever.
Code generation with planning: Chain-of-Thought (CoT) is a popular strategy in code generation [60, 25, 20]. For instance, CodeCoT [20] leverages CoT to break down the requirements into steps described in natural language and then convert them to code. Some other works employ planning strategies [42, 2, 49, 61]. As an example, PairCoder [61] incorporates two collaborative LLM agents. The navigator agent proposes solution plans, selects the optimal plan, and controls the next iteration round based on execution feedback. The driver agent follows its guidance and performs initial code generation, testing, and refinement.
Code generation with iterative refinement: Multi-agent systems can dynamically refine the generated code through iterative feedback from models [53, 27, 4, 21], tools [62, 56, 63, 62, 17, 53, 42, 50, 10], and humans [51, 35]. For instance, the agent can receive peer-reflection from multiple models with specialized roles [53, 27] or conduct self-refinement based on the previous output [4, 21]. The agent can also collect informative compilation errors from compiler and interpreter tools [62, 56], or even get access to external resources by applying retrieval or searching tools [63, 62, 17]. Furthermore, hybrid feedback from tool and model is possible – receiving error messages from tools and then using LLMs’ contextual understanding to provide corresponding explanations, suggestions, and instructions [53, 42, 50]. Considering the critical role of human in clarifying ambiguous requirements, several works incorporate human feedback into the code generation process [51, 35].
3.3 Static Code Checking
Static code checking refers to examining the quality of code without executing it, which is an essential step to identify diverse code quality issues, such as bugs, vulnerabilities, or code smells before executing the tests extensively.
Static bug detection: Preliminary studies show that fine-tuning LLMs on code base or simply prompting them has demonstrated superior effectiveness in identifying bugs in code snippets [12, 58]. From the perspective of multi-agent collaboration, an approach for vulnerability detection [33] simulates the real-world code review process with multiple test and developer agents. GPTLENS [19] creates a synergistic framework, where several auditor agents randomly generate potential vulnerabilities and their corresponding reasoning thoroughly, while the critic agents carefully review and score them based on specific criteria. LLIFT [28] combines LLM-based agentic systems with traditional static analysis tools. Based on the undecided bugs reported by the tool UBITect, LLIFT further leverages LLMs’ capability in code comprehension and summarization to identify the UBI bugs in Linux kernels.
Code review: In code review, LLM-based agents mimic the real-world peer review procedure by including multiple agents as different code reviewers. CodeAgent [44] simulates a waterfall-like pipeline with four stages and sets up a code review team with six agents of different characters, including user, CEO, CPO, CTO, coder, and reviewer. ICAA [11] designs a multi-agent system to identify code-intention inconsistencies, where the context incubation agent gathers essential information from code repositories, the consistency checking agent analyzes this information to identify any discrepancies, and the report agent compiles the final report.
3.4 Testing and Quality Assurance
Creating tests in practice can be challenging, as test cases directly generated by LLMs can encounter compilation or execution errors. The tests must be both syntactically and semantically accurate, while also being sufficiently comprehensive to cover as many different states of the software under test as possible.
Unit test: LLM-based agents can eliminate unit test errors by iteratively collecting error messages and fixing the buggy test code [59, 41]. Similarly, CoverUp [37] refines its prompts iteratively to focus on the code snippets that lack coverage, thus achieving higher test coverage rates.
System test: WhiteFox [55] encompasses two agents working together. An analysis agent examines the low-level source code and produces requirements on the high-level test programs that can trigger the optimization, while a generation agent crafts test programs based on summarized requirements. In terms of mobile applications, GPTDroid [30] automates the entire process by perceiving GUI page information, generating test scripts in the form of Q&A, executing these scripts using tools, and receiving feedback from the application. Regarding web applications, given an API name, RESTSpecIT [6] generates and mutates HTTP requests through a reflection loop. By sending these requests to the API endpoint, it analyzes the HTTP responses for inference and testing. The agent then uses valid requests as feedback to refine the mutations in each iteration.
3.5 Debugging
Fault localization: AgentFL [38] scales up LLM-based fault localization to project-level code context through the synergy of multiple agents. AUTOFL [23], a single-agent system, enhances standalone LLMs with tool invocation, utilizing four specialized function calls to explore the repository more effectively.
Program repair: Feedback-driven approaches dominate recent program repair research. ChatRepair[54] refines patches based on iterative environmental feedback, while CigaR[18] applies similar principles to function-level repair. AutoSD [24] adopts a scientific debugging methodology, integrating hypothesis generation, execution-based validation, conclusion making, and automated fixing into a unified repair pipeline.
4 Methodological Approaches and Research Agenda
4.1 Model Selection and Evaluation
The effectiveness of multi-agent systems fundamentally depends on the capabilities of the underlying language models. However, deploying such systems in real-world software engineering scenarios requires careful tradeoffs between model performance, computational cost, and data privacy.
Proprietary frontier models represent performance upper bounds. Notable examples include GPT-5 [36] from OpenAI, Gemini-3-pro [15] from Google, and Claude-v4.5 [1] developed by Anthropic, which demonstrate forefront performance in language understanding and generation. However, these models require API-based access, which raises significant privacy and security concerns when processing proprietary codebases.
Open-source alternatives address privacy and cost constraints while enabling local deployment and customization. For instance, LLaMA-4 [34] and DeepSeek-V3 [29] can achieve competitive performance to proprietary models. Such open-source LLMs enable organizations to maintain full control over sensitive codebases, eliminate recurring API costs, and customize model behavior for domain-specific SE tasks, though requiring higher infrastructure investment.
Reasoning-optimized architectures represent a recent paradigm shift in model design. Compared with contemporaneous models, DeepSeek-R1 incorporates specialized reasoning capabilities, which yields advantages for complex software engineering tasks like debugging and architectural decision-making. As new competent models emerge, evaluating them systematically across various SE tasks allows us to discover alternatives that balance cost-efficiency, privacy protection, and strong performance in addressing complex software engineering problems.
4.2 Comparative Analysis and Benchmarks
The software engineering domain offers a rich ecosystem of open-source datasets and benchmarks spanning diverse tasks. For instance in requirements engineering, PURE [13] and PROMISE [26] datasets provide standardized evaluation data for software requirement classification. For code generation, humanEval [3] emphasizes the functional correctness on programming problems, while GSM8K [5] assesses mathematical reasoning in code. With respect to quality assurance, BugBench [31] provides a benchmark for bug detection. Tests4Py [43] focuses on system testing scenarios. By evaluating LLM-based solutions across these established benchmarks, we provide empirical evidence for the effectiveness of our approaches while enabling direct comparison with existing methods.
4.3 Agentic Frameworks and Coordination
To facilitate the development and deployment of agentic solutions, the community has proposed a wide range of frameworks for agents, including but not limited to CrewAI, LangGraph, Google ADK, AutoGen, and n8n. The rapid growth of these frameworks highlights the need to evaluate their architectural principles, communication mechanisms, memory management strategies, and safety guardrails [7].
In parallel with these frameworks, agent communication protocols have been introduced to standardize inter-agent interactions and promote collaborative problem-solving [57]. Widely-used examples include Contract Net Protocol (CNP), Agent-to-Agent (A2A), and Agent Network Protocol (ANP). Meanwhile, Model Context Protocol (MCP) has emerged as an open standard that unifies the integration of external tools and data sources.
Despite the abundance and diversity of these frameworks and protocols, deploying them in real-world software projects raises significant concerns on reliability and privacy. As development teams transition from proof-of-concept demonstrations to production environments, they might encounter critical challenges, including dynamic agent discovery, seamless integration of human feedback, and effective management of computational costs and data constraints. These practical deployment gaps motivate the investigation of key challenges and research opportunities in Section 5.
5 Future Challenges and Opportunities
Despite the promising future of LLM-based multi-agent systems throughout the Software Development Life Cycle (SDLC), numerous challenges remain:
  • Enhancing individual agents’ capabilities: Current LLM-based agentic systems can only simulate generic roles like software developers and product managers. They still lack the nuanced expertise to mimic more SE-specific roles, such as for vulnerability detection or security auditing. This shortcoming emphasizes the need to integrate domain-specific knowledge into LLMs to support specialized SE roles. Meanwhile, integrating effective prompting frameworks can significantly facilitate agentic systems. Investigating more efficient Retrieval Augmented Generation (RAG) [14] techniques and agent-based workflows will remain a prominent area of research in the future.
  • Optimizing agent synergy and human-agent coordination: Given the diverse strength of individual agents, the next phase of research focuses on improving the internal dynamics among agents and the external human interaction with multi-agent systems. While some studies have incorporated human-in-the-loop designs, e.g., MARE system [22] for requirement engineering, many research questions remain unanswered. These include optimizing human roles, enhancing feedback mechanisms, and identifying appropriate intervention points in human-agent collaboration.
  • Collecting more data throughout the SDLC: Most agents rely on general-purpose LLMs such as ChatGPT or code-specific models like Deepseek-Coder. However, complex SE tasks require more specific data beyond code. Collecting and utilizing valuable data from the entire SDLC, such as design, architecture, developer discussions, historical code changes, and dynamic runtime information, can enhance LLMs’ capabilities significantly.
  • Reducing computational efforts and costs: Reducing the costs associated with deploying and maintaining LLM-based multi-agent systems is crucial for ensuring scalability, sustainability, and cost-effectiveness in real-world projects. Distilling large-scale models or fine-tuning light-weight models for a specific SE agent could be a solution; incorporating robust feedback mechanisms and minimizing iterations between agents can also lower the computational costs for LLM inference.
  • Evaluating multi-agent systems for SE: Existing benchmarks have focused on individual SE tasks and exhibit limitations as SE projects grow in complexity. However, software engineering is inherently collaborative, including activities like joint requirements gathering, code integration, and peer reviews. Current benchmarks often overlook these aspects, highlighting an increasing need to assess LLMs’ cooperative capabilities within multi-agent settings.
6 Conclusion
In this work, we reviewed the most recent advancements in LLM-based multi-agent systems for SE, proposed potential research directions, and discussed future challenges for practical deployment. Our analysis reveals their promise across various SDLC stages, but also emphasizes that fully automated software development requires progress beyond individual stage. Software engineering is inherently collaborative, involving collective decision-making, iterative peer refinement, and continuous coordination among diverse roles.
As for future work, We plan to compare multi-agent design patterns, such as planner–executor pairs, self-refinement loops, and human-in-the-loop, and perform empirical evaluations on more open-source and proprietary LLMs. To move forward and exploit the full potential of agentic systems for SE, we encourage the entire community to work jointly and overcome current limitations together.
Acknowledgments
This work was carried out within the ITEA 4 project GENIUS, as part of the ITEA programme, the Eureka Cluster on software innovation. This work was funded by the German Federal Ministry of Research, Technology and Space (BMFTR).
Literatur
  • [1] Anthropic (2025) Introducing Claude Sonnet 4.5d. Note: Accessed September 29 External Links: Link Cited by: §4.1.
  • [2] R. Bairi, A. Sonwane, A. Kanade, A. Iyer, S. Parthasarathy, S. Rajamani, B. Ashok, and S. Shet (2024) Codeplan: repository-level coding using llms and planning. Proceedings of the ACM on Software Engineering 1 (FSE), pp. 675–698. Cited by: §3.2.
  • [3] M. Chen, J. Tworek, H. Jun, Q. Yuan, H. P. de Oliveira Pinto, J. Kaplan, H. Edwards, Y. Burda, N. Joseph, et al. (2021) Evaluating large language models trained on code. External Links: 2107.03374 Cited by: §4.2.
  • [4] X. Chen, M. Lin, N. Schärli, and D. Zhou (2023) Teaching large language models to self-debug. arXiv preprint arXiv:2304.05128. Cited by: §3.2.
  • [5] K. Cobbe, V. Kosaraju, M. Bavarian, M. Chen, H. Jun, L. Kaiser, M. Plappert, J. Tworek, J. Hilton, R. Nakano, C. Hesse, and J. Schulman (2021) Training verifiers to solve math word problems. arXiv preprint arXiv:2110.14168. Cited by: §4.2.
  • [6] A. Decrop, G. Perrouin, M. Papadakis, X. Devroey, and P. Schobbens (2024) You can rest now: automated specification inference and black-box testing of restful apis with large language models. arXiv preprint arXiv:2402.05102. Cited by: §3.4.
  • [7] H. Derouiche, Z. Brahmi, and H. Mazeni (2025) Agentic ai frameworks: architectures, protocols, and design challenges. External Links: 2508.10146, Link Cited by: §4.3.
  • [8] J. Devlin, M. Chang, K. Lee, and K. Toutanova (2019) BERT: pre-training of deep bidirectional transformers for language understanding. In North American Chapter of the Association for Computational Linguistics, External Links: Link Cited by: §2.1.
  • [9] Q. Dong, L. Li, D. Dai, C. Zheng, J. Ma, R. Li, H. Xia, J. Xu, Z. Wu, B. Chang, et al. (2024) A survey on in-context learning. In Proceedings of the 2024 conference on empirical methods in natural language processing, pp. 1107–1128. Cited by: §1.
  • [10] H. Elhashemy, Y. Lotfy, and Y. Tang (2025) Bridging the prototype-production gap: a multi-agent system for notebooks transformation. arXiv preprint arXiv:2511.07257. Cited by: §3.2.
  • [11] G. Fan, X. Xie, X. Zheng, Y. Liang, and P. Di (2023) Static code analysis in the ai era: an in-depth exploration of the concept, function, and potential of intelligent code analysis agents. arXiv preprint arXiv:2310.08837. Cited by: §3.3.
  • [12] S. Feng and C. Chen (2024) Prompting is all you need: automated android bug replay with large language models. In Proceedings of the 46th IEEE/ACM International Conference on Software Engineering, pp. 1–13. Cited by: §3.3.
  • [13] A. Ferrari, G. O. Spagnolo, and S. Gnesi (2017) Pure: a dataset of public requirements documents. In 2017 IEEE 25th international requirements engineering conference (RE), pp. 502–505. Cited by: §4.2.
  • [14] Y. Gao, Y. Xiong, X. Gao, K. Jia, J. Pan, Y. Bi, J. Sun, and H. Wang (2023) Retrieval-augmented generation for large language models: a survey. Cited by: §1, 1st item.
  • [15] Google (2025) Gemini 3 pro - best for complex tasks and bringing creative concepts to life. External Links: Link Cited by: §4.1.
  • [16] Z. Han, C. Gao, J. Liu, J. Zhang, and S. Q. Zhang (2024) Parameter-efficient fine-tuning for large models: a comprehensive survey. Transactions on Machine Learning Research. Cited by: §1.
  • [17] X. He, J. Zou, Y. Lin, M. Zhou, S. Han, Z. Yuan, and D. Zhang (2024) CoCoST: automatic complex code generation with online searching and correctness testing. arXiv preprint arXiv:2403.13583. Cited by: §3.2.
  • [18] D. Hidvégi, K. Etemadi, S. Bobadilla, and M. Monperrus (2024) Cigar: cost-efficient program repair with llms. arXiv preprint arXiv:2402.06598. Cited by: §3.5.
  • [19] S. Hu, T. Huang, F. İlhan, S. F. Tekin, and L. Liu (2023) Large language model-powered smart contract vulnerability detection: new perspectives. In 2023 5th IEEE International Conference on Trust, Privacy and Security in Intelligent Systems and Applications (TPS-ISA), pp. 297–306. Cited by: §3.3.
  • [20] D. Huang, Q. Bu, Y. Qing, and H. Cui (2023) Codecot: tackling code syntax errors in cot reasoning for code generation. CoRR 2308, pp. 1–20. Cited by: §3.2.
  • [21] S. Jiang, Y. Wang, and Y. Wang (2023) Selfevolve: a code evolution framework via large language models. arXiv preprint arXiv:2306.02907. Cited by: §3.2.
  • [22] D. Jin, Z. Jin, X. Chen, and C. Wang (2024) MARE: multi-agents collaboration framework for requirements engineering. arXiv preprint arXiv:2405.03256. Cited by: §3.1, 2nd item.
  • [23] S. Kang, G. An, and S. Yoo (2023) A preliminary evaluation of llm-based fault localization. arXiv preprint arXiv:2308.05487. Cited by: §3.5.
  • [24] S. Kang, B. Chen, S. Yoo, and J. Lou (2025) Explainable automated debugging via large language model-driven scientific debugging. Empirical Software Engineering 30 (2), pp. 1–28. Cited by: §3.5.
  • [25] H. Le, H. Chen, A. Saha, A. Gokul, D. Sahoo, and S. Joty (2023) Codechain: towards modular code generation through chain of self-revisions with representative sub-modules. arXiv preprint arXiv:2310.08992. Cited by: §3.2.
  • [26] F. Li, J. Horkoff, J. Mylopoulos, R. S. Guizzardi, G. Guizzardi, A. Borgida, and L. Liu (2014) Non-functional requirements as qualities, with a spice of ontology. In 2014 IEEE 22nd International Requirements Engineering Conference (RE), pp. 293–302. Cited by: §4.2.
  • [27] G. Li, H. Hammoud, H. Itani, D. Khizbullin, and B. Ghanem (2023) Camel: communicative agents for" mind" exploration of large language model society. Advances in Neural Information Processing Systems 36, pp. 51991–52008. Cited by: §3.2.
  • [28] H. Li, Y. Hao, Y. Zhai, and Z. Qian (2024) Enhancing static analysis for practical bug detection: an llm-integrated approach. Proceedings of the ACM on Programming Languages 8 (OOPSLA1), pp. 474–499. Cited by: §3.3.
  • [29] A. Liu, B. Feng, B. Xue, B. Wang, B. Wu, C. Lu, C. Zhao, C. Deng, C. Zhang, C. Ruan, et al. (2024) Deepseek-v3 technical report. arXiv preprint arXiv:2412.19437. Cited by: §4.1.
  • [30] Z. Liu, C. Chen, J. Wang, M. Chen, B. Wu, X. Che, D. Wang, and Q. Wang (2024) Make llm a testing expert: bringing human-like interaction to mobile gui testing via functionality-aware decisions. In Proceedings of the IEEE/ACM 46th International Conference on Software Engineering, pp. 1–13. Cited by: §3.4.
  • [31] S. Lu, Z. Li, F. Qin, L. Tan, P. Zhou, and Y. Zhou (2005) Bugbench: benchmarks for evaluating bug detection tools. In Workshop on the evaluation of software defect detection tools, Vol. 5. Cited by: §4.2.
  • [32] L. Ma, S. Liu, Y. Li, X. Xie, and L. Bu (2024) SpecGen: automated generation of formal program specifications via large language models. arXiv preprint arXiv:2401.08807. Cited by: §3.1.
  • [33] Z. Mao, J. Li, D. Jin, M. Li, and K. Tei (2024) Multi-role consensus through llms discussions for vulnerability detection. In 2024 IEEE 24th International Conference on Software Quality, Reliability, and Security Companion (QRS-C), pp. 1318–1319. Cited by: §3.3.
  • [34] Meta (2025) LLaMA 4: leading multimodal intelligence. External Links: Link Cited by: §4.1.
  • [35] F. Mu, L. Shi, S. Wang, Z. Yu, B. Zhang, C. Wang, S. Liu, and Q. Wang (2023) ClarifyGPT: empowering llm-based code generation with intention clarification. arXiv preprint arXiv:2310.10996. Cited by: §3.2.
  • [36] OpenAI (2025) GPT-5 system card. Note: Accessed August 7 External Links: Link Cited by: §4.1.
  • [37] J. A. Pizzorno and E. D. Berger (2024) CoverUp: coverage-guided llm-based test generation. arXiv preprint arXiv:2403.16218. Cited by: §3.4.
  • [38] Y. Qin, S. Wang, Y. Lou, J. Dong, K. Wang, X. Li, and X. Mao (2024) AgentFL: scaling llm-based fault localization to project-level context. arXiv preprint arXiv:2403.16362. Cited by: §3.5.
  • [39] P. Sahoo, A. K. Singh, S. Saha, V. Jain, S. Mondal, and A. Chadha (2024) A systematic survey of prompt engineering in large language models: techniques and applications. arXiv preprint arXiv:2402.07927. Cited by: §1.
  • [40] M. A. Sami, M. Waseem, Z. Zhang, Z. Rasheed, K. Systä, and P. Abrahamsson (2024) Early results of an ai multiagent system for requirements elicitation and analysis. In International Conference on Product-Focused Software Process Improvement, pp. 307–316. Cited by: §3.1.
  • [41] M. Schäfer, S. Nadi, A. Eghbali, and F. Tip (2023) An empirical evaluation of using large language models for automated unit test generation. IEEE Transactions on Software Engineering. Cited by: §3.4.
  • [42] N. Shinn, F. Cassano, A. Gopinath, K. Narasimhan, and S. Yao (2024) Reflexion: language agents with verbal reinforcement learning. Advances in Neural Information Processing Systems 36. Cited by: §3.2, §3.2.
  • [43] M. Smytzek, M. Eberlein, B. Serce, L. Grunske, and A. Zeller (2024) Tests4Py: a benchmark for system testing. In Companion Proceedings of the 32nd ACM International Conference on the Foundations of Software Engineering, pp. 557–561. Cited by: §4.2.
  • [44] D. Tang, Z. Chen, K. Kim, Y. Song, H. Tian, S. Ezzini, Y. Huang, J. Klein, and T. F. Bissyandé (2024) CodeAgent: collaborative agents for software engineering. ArXiv abs/2402.02172. External Links: Link Cited by: §3.3.
  • [45] Y. Tang, R. Hasan, and T. Runkler (2024) FsPONER: Few-shot prompt optimization for named entity recognition in domain-specific scenarios. In ECAI 2024, pp. 3757–3764. External Links: Link Cited by: §1.
  • [46] Y. Tang, D. Tuncel, C. Koerner, and T. Runkler (2025) The few-shot dilemma: over-prompting large language models. arXiv preprint arXiv:2509.13196. Cited by: §1.
  • [47] N. N. Vaidya, T. A. Runkler, T. Hubauer, V. Haderlein-Hoegberg, and M. M. Brandt (2025) Conceptual in-context learning and chain of concepts: solving complex conceptual problems using large language models. In 2025 IEEE Symposium on Computational Intelligence in Natural Language Processing and Social Media (CI-NLPSoMe), Vol. , pp. 1–7. External Links: Document Cited by: §1.
  • [48] A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, Ł. Kaiser, and I. Polosukhin (2017) Attention is all you need. Advances in Neural Information Processing Systems 30. Cited by: §2.1.
  • [49] C. Wang, J. Zhang, Y. Feng, T. Li, W. Sun, Y. Liu, and X. Peng (2024) Teaching code llms to use autocompletion tools in repository-level code generation. arXiv preprint arXiv:2401.06391. Cited by: §3.2.
  • [50] H. Wang, Z. Liu, S. Wang, G. Cui, N. Ding, Z. Liu, and G. Yu (2024) Intervenor: prompting the coding ability of large language models with the interactive chain of repair. In Findings of the Association for Computational Linguistics ACL 2024, pp. 2081–2107. Cited by: §3.2.
  • [51] X. Wang, Z. Wang, J. Liu, Y. Chen, L. Yuan, H. Peng, and H. Ji (2023) Mint: evaluating llms in multi-turn interaction with tools and language feedback. arXiv preprint arXiv:2309.10691. Cited by: §3.2.
  • [52] J. Wei, Y. Tay, R. Bommasani, C. Raffel, B. Zoph, S. Borgeaud, D. Yogatama, M. Bosma, D. Zhou, D. Metzler, E. H. Chi, T. Hashimoto, O. Vinyals, P. Liang, J. Dean, and W. Fedus (2022) Emergent abilities of large language models. External Links: 2206.07682 Cited by: §2.1.
  • [53] Q. Wu, G. Bansal, J. Zhang, Y. Wu, B. Li, E. Zhu, L. Jiang, X. Zhang, S. Zhang, J. Liu, et al. (2024) Autogen: enabling next-gen llm applications via multi-agent conversations. In First Conference on Language Modeling, Cited by: §3.2.
  • [54] C. S. Xia and L. Zhang (2024) Automated program repair via conversation: fixing 162 out of 337 bugs for \$0.42 each using chatgpt. In Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis, pp. 819–831. Cited by: §3.5.
  • [55] C. Yang, Y. Deng, R. Lu, J. Yao, J. Liu, R. Jabbarvand, and L. Zhang (2024) Whitefox: white-box compiler fuzzing empowered by large language models. Proceedings of the ACM on Programming Languages 8 (OOPSLA2), pp. 709–735. Cited by: §3.4.
  • [56] J. Yang, A. Prabhakar, K. Narasimhan, and S. Yao (2024) Intercode: standardizing and benchmarking interactive coding with execution feedback. Advances in Neural Information Processing Systems 36. Cited by: §3.2.
  • [57] Y. Yang, H. Chai, Y. Song, S. Qi, M. Wen, N. Li, J. Liao, H. Hu, J. Lin, G. Chang, et al. (2025) A survey of ai agent protocols. arXiv preprint arXiv:2504.16736. Cited by: §4.3.
  • [58] Z. Yuan, J. Liu, Q. Zi, M. Liu, X. Peng, and Y. Lou (2023) Evaluating instruction-tuned large language models on code comprehension and generation. arXiv preprint arXiv:2308.01240. Cited by: §3.3.
  • [59] Z. Yuan, Y. Lou, M. Liu, S. Ding, K. Wang, Y. Chen, and X. Peng (2023) No more manual tests? evaluating and improving chatgpt for unit test generation. arXiv preprint arXiv:2305.04207. Cited by: §3.4.
  • [60] E. Zelikman, Q. Huang, G. Poesia, N. Goodman, and N. Haber (2023) Parsel: algorithmic reasoning with language models by composing decompositions. Advances in Neural Information Processing Systems 36, pp. 31466–31523. Cited by: §3.2.
  • [61] H. Zhang, W. Cheng, Y. Wu, and W. Hu (2024) A pair programming framework for code generation via multi-plan exploration and feedback-driven refinement. In Proceedings of the 39th IEEE/ACM International Conference on Automated Software Engineering, pp. 1319–1331. Cited by: §3.2.
  • [62] K. Zhang, J. Li, G. Li, X. Shi, and Z. Jin (2024) Codeagent: enhancing code generation with tool-integrated agent systems for real-world repo-level coding challenges. arXiv preprint arXiv:2401.07339. Cited by: §3.2.
  • [63] K. Zhang, H. Zhang, G. Li, J. Li, Z. Li, and Z. Jin (2023) Toolcoder: teach code generation models to use api search tools. arXiv preprint arXiv:2305.04032. Cited by: §3.2.
  • [64] W. X. Zhao, K. Zhou, J. Li, T. Tang, X. Wang, Y. Hou, Y. Min, B. Zhang, J. Zhang, Z. Dong, Y. Du, C. Yang, Y. Chen, Z. Chen, J. Jiang, R. Ren, Y. Li, X. Tang, Z. Liu, P. Liu, J. Nie, and J. Wen (2024) A survey of large language models. External Links: 2303.18223, Link Cited by: §1.